### v1

# Disturbance-Robust Backup Control Barrier Functions: Safety Under Uncertain Dynamics

Abstract
Obtaining a controlled invariant set is crucial for safety-critical control with control barrier functions (CBFs) but is non-trivial for complex nonlinear systems and constraints. Backup control barrier functions allow such sets to be constructed online in a computationally tractable manner by examining the evolution (or flow) of the system under a known backup control law. However, for systems with unmodeled disturbances, this flow cannot be directly computed, making the current methods inadequate for assuring safety in these scenarios. To address this gap, we leverage bounds on the nominal and disturbed flow to compute a forward invariant set online by ensuring safety of an expanding norm ball tube centered around the nominal system evolution. We prove that this set results in robust control constraints which guarantee safety of the disturbed system via our Disturbance-Robust Backup Control Barrier Function (DR-BCBF) solution. Additionally, the efficacy of the proposed framework is demonstrated in simulation, applied to a double integrator problem and a rigid body spacecraft rotation problem with rate constraints.
I Introduction
Control barrier functions (CBFs) [1, 2], are a popular approach to assuring safety of autonomous systems by encoding notions of safety directly into existing controllers and providing sufficient conditions for forward invariance of safe sets. However, obtaining safe sets for which every point has a safe control action (i.e., is controlled invariant) is difficult for high-dimensional systems. Additionally, dynamics models are seldom perfect. In this paper we seek to solve both of these problems simultaneously.
To address the problem of controlled invariance, we adapt the backup set method from [3, 4] based on online backward reachability. This method establishes a controlled invariant safe set implicitly using the flow of the system under a prescribed backup control law. This approach is computationally tractable even for complex systems. For affine nonlinear systems, this technique generates linear control constraints which can be used to efficiently solve for point-wise optimal control signals for an arbitrary primary controller.
The second problem we address is that of model uncertainty, in the form of unknown bounded disturbances. Much work has been done in recent years in the CBF literature to address model uncertainty and disturbances. In [5] robust CBFs (RCBFs) are introduced which account for the worst-case disturbance. RCBFs with input constraints were considered in [6] with an application to spacecraft control. In [7] the authors synthesized RCBFs and control Lyapunov functions together and provided additional conditions for convergence guarantees. RCBFs for relative degree two, sampled-data systems were developed and applied to robotic grasping in [8]. However, these approaches all assume that an explicit representation of a controlled invariant safe set can be found. Unlike these approaches, [9] leverages the mixed-monotonicity property of certain systems to over-approximate the reachable sets computed using the backup set method.
Figure 1: Overview of the proposed disturbance-robust safety-critical control framework. $\mathcal{C}_{\rm I}$ represents a forward invariant subset of an unknown controlled invariant set $\mathcal{C}_{\rm D}$ and guarantees safety of the disturbed system.
This manuscript presents a robust safety-critical control framework, displayed in Figure 1, which extends the method of backup CBFs to make formal safety guarantees on systems with unknown disturbances. The main contributions of this paper are two-fold. First, we introduce disturbance-robust backup CBFs from which we derive tractable robust safety conditions which can be synthesized directly into existing controllers. Second, we demonstrate the effectiveness of the approach using two numerical simulations: a toy problem and a spacecraft rotation example.
II Preliminaries
This section reviews the fundamental concepts of safety-critical control using control barrier functions (CBFs) and constructing invariant sets using an implicit approach, also referred to as the backup set method.
II-A Control Barrier Functions
Consider a nonlinear control affine dynamical system modeled as
$\dot{\boldsymbol{x}}=f(\boldsymbol{x})+g(\boldsymbol{x})\boldsymbol{u},$ (1)
where $f:\mathcal{X}\rightarrow\mathbb{R}^{n}$ and $g:\mathcal{X}\rightarrow\mathbb{R}^{n\times m}$ are Lipschitz continuous functions, $\boldsymbol{x}\in\mathcal{X}\subseteq\mathbb{R}^{n}$ represents the state vector, and $\boldsymbol{u}\in\mathcal{U}\subseteq\mathbb{R}^{m}$ represents the control vector. $\mathcal{X}$ is the set of all possible states and $\mathcal{U}$ is the admissible control set. For an initial condition $\boldsymbol{x}(0)=\boldsymbol{x}_{0}\in\mathcal{X}$ if $\boldsymbol{u}$ is given by a locally Lipschitz feedback controller $k:\mathcal{X}\rightarrow\mathcal{U}$, $\boldsymbol{u}=k(\boldsymbol{x})$, the closed-loop system (1) has a unique solution $\boldsymbol{\phi}^{n}(t,\boldsymbol{x}_{0})$ over an interval of existence.
In the context of this work, safety is defined as a region of the state space, known as the safe set $\mathcal{C}_{\rm S}$, within which the dynamical system must stay. Safe controllers are ones that render this safe set forward invariant. A set $\mathcal{C}\subset\mathbb{R}^{n}$ is forward invariant along (1) if $\boldsymbol{x}_{0}\in\mathcal{C}\implies\boldsymbol{\phi}^{n}(t,\boldsymbol{x}_{0})\in\mathcal{C},$ for all $t>0$. Now, consider the safe set $\mathcal{C}_{\rm S}$ as the 0-superlevel set of a continuously differentiable function $h:\mathcal{X}\rightarrow\mathbb{R}$ with
$\mathcal{C}_{\rm S}\triangleq\{\boldsymbol{x}\in\mathcal{X}:h(\boldsymbol{x})\geq 0\},$ (2)
and the gradient along the boundary of $\mathcal{C}_{\rm S}$ remains nonzero, i.e., $h(\boldsymbol{x})=0\implies\nabla h(\boldsymbol{x})\neq 0$.
Definition 1.
[…]
$\sup_{\boldsymbol{u}\in\mathcal{U}}\dot{h}(\boldsymbol{x},\boldsymbol{u})\triangleq\underbrace{\nabla h(\boldsymbol{x})f(\boldsymbol{x})}_{L_{f}h(\boldsymbol{x})}+\underbrace{\nabla h(\boldsymbol{x})g(\boldsymbol{x})}_{L_{g}h(\boldsymbol{x})}\boldsymbol{u}\geq-\alpha(h(\boldsymbol{x})),$
where $L_{f}h$ and $L_{g}h$ are the Lie derivatives of $h$ along $f$ and $g$, respectively.
Finally, we arrive at the main result of [1], relating CBFs and forward invariance to safety.
Theorem 1.
If $h$ is a CBF for (1) on $\mathcal{C}_{\rm S}$, then any locally Lipschitz controller $k:\mathcal{X}\rightarrow\mathcal{U}$, $\boldsymbol{u}=k(\boldsymbol{x})$ satisfying
$L_{f}h(\boldsymbol{x})+L_{g}h(\boldsymbol{x})\boldsymbol{u}\geq-\alpha(h(\boldsymbol{x}))$ (3)
for all $\boldsymbol{x}\in\mathcal{C}_{\rm S}$ will render the set $\mathcal{C}_{\rm S}$ forward invariant.
The sufficient condition for forward invariance of the safe set is linear in the control, often motivating the use of CBFs to supply safety constraints to a quadratic program (QP). For an arbitrary primary controller $\boldsymbol{u}_{\rm p}\in\mathcal{U}$, it is possible to ensure safety of (1) by solving the following point-wise optimization problem for the safe control, $\boldsymbol{u}_{\rm safe}$:
$\begin{gathered}\boldsymbol{u}_{\rm safe}=\underset{\boldsymbol{u}\in\mathcal{U}}{\text{argmin}}\mkern 9.0mu\frac{1}{2}\left\|\boldsymbol{u}_{\rm p}-\boldsymbol{u}\right\|^{2}\quad\quad\quad\quad(\text{CBF-QP})\\ \quad\quad\quad\quad\text{s.t.}\quad L_{f}h(\boldsymbol{x})+L_{g}h(\boldsymbol{x})\boldsymbol{u}\geq-\alpha(h(\boldsymbol{x})).\end{gathered}$
A key challenge is actually obtaining an explicit representation of such a function $h$ where a safe control signal satisfying (3) can always be found. Depending on the safe set this may be difficult or impossible, especially for high dimensional systems. This therefore motivates the use of an extension of CBFs known as backup CBFs.
II-B Implicitly Defined Invariant Sets
First introduced in [3] and expanded upon in [4], the backup set approach relies on filtering with respect to an implicitly defined controlled invariant set.
Definition 2.
A set $\mathcal{C}\subset\mathbb{R}^{n}$ is controlled invariant if there exists a controller $k:\mathcal{X}\rightarrow\mathcal{U}$, $\boldsymbol{u}=k(\boldsymbol{x})$ which renders $\mathcal{C}$ forward invariant for a dynamical system defined by (1).
To construct an implicit controlled invariant set, first assume that we have defined a set $\mathcal{C}_{\rm S}$ describing our state constraints, which is not necessarily controlled invariant. Now suppose that there exists a set within $\mathcal{C}_{\rm S}$ which we call a backup set, $\mathcal{C}_{\rm B}$, such that $\mathcal{C}_{\rm B}\subset\mathcal{C}_{\rm S}$. This set is defined similar to (2) with a continuously differentiable function $h_{\rm b}$, it is known to be controlled invariant, and it is made forward invariant by a backup control law defined by $\boldsymbol{u}_{\rm b}:\mathcal{X}\rightarrow\mathcal{U}$. This backup control law and set can often be obtained via an application of Lyapunov stability theory around local regions of attraction [10]. The closed-loop system under such a backup control law is denoted
$f_{\rm cl}(\boldsymbol{x})\triangleq f(\boldsymbol{x})+g(\boldsymbol{x})\boldsymbol{u}_{\rm b}(\boldsymbol{x}).$ (4)
To obtain this implicitly defined controlled invariant set, $\mathcal{C}_{\rm BI}$, satisfying $\mathcal{C}_{\rm B}\subseteq\mathcal{C}_{\rm BI}\subseteq\mathcal{C}_{\rm S}$, one must ensure that the trajectory of the system under $\boldsymbol{u}_{\rm b}(\boldsymbol{x})$ remains in $\mathcal{C}_{\rm S}$ over a finite time horizon, and that the final point in the trajectory lies within the backup set $\mathcal{C}_{\rm B}$. It is assumed that for any $\boldsymbol{x}\in\mathcal{X}$ there exists a unique solution $\boldsymbol{\phi}^{n}_{\rm b}:[0,T]\times\mathcal{X}\rightarrow\mathcal{X}$ which satisfies:
${\dot{\boldsymbol{\phi}^{n}_{\rm b}\mkern-3.0mu}\mkern 3.0mu}{}(\tau,\boldsymbol{x})=f_{\rm cl}(\boldsymbol{\phi}^{n}_{\rm b}(\tau,\boldsymbol{x})),\mkern 9.0mu\boldsymbol{\phi}^{n}_{\rm b}(0,\boldsymbol{x})=\boldsymbol{x}.$ (5)
The solution is the flow of the system over the interval $[0,T]$ for $T\in\mathbb{R}_{>0}$ starting at state $\boldsymbol{x}$ under the backup control law $\boldsymbol{u}_{\rm b}$. Therefore $\mathcal{C}_{\rm BI}$ is defined as
$\mathcal{C}_{\rm BI}\triangleq\left\{\boldsymbol{x}\in\mathcal{X}\,\middle|\,\begin{array}[]{c}h(\boldsymbol{\phi}^{n}_{\rm b}(\tau,\boldsymbol{x}))\geq 0,\forall\mkern 1.0mu\tau\in[0,T],\\ h_{\rm b}(\boldsymbol{\phi}^{n}_{\rm b}(T,\boldsymbol{x}))\geq 0\\ \end{array}\right\}.$ (8)
While $\mathcal{C}_{\rm BI}$ is seldom the largest controlled invariant subset of $\mathcal{C}_{\rm S}$ (known as the viability kernel [11]), it is an under-approximation of such a set which can be computed in real-time.
As before, these nonlinear inequalities on the state which define the controlled invariant set $\mathcal{C}_{\rm BI}$ can be converted into linear control constraints. We require $\forall\mkern 2.0mu\tau\in[0,T]$, and class-$\mathcal{K}_{\infty}$ functions $\alpha$ and $\alpha_{\rm b}$ that
$\dot{h}(\boldsymbol{\phi}^{n}_{\rm b}(\tau,\boldsymbol{x}),\boldsymbol{u})$ $\geq-\alpha(h(\boldsymbol{\phi}^{n}_{\rm b}(\tau,\boldsymbol{x}))),$ (9a)
$\dot{h}_{\rm b}(\boldsymbol{\phi}^{n}_{\rm b}(T,\boldsymbol{x}),\boldsymbol{u})$ $\geq-\alpha_{\rm b}(h_{\rm b}(\boldsymbol{\phi}^{n}_{\rm b}(T,\boldsymbol{x}))).$ (9b)
Expanding the total derivative terms, the constraints can be written as
$\nabla h(\boldsymbol{\phi}^{n}_{\rm b}(\tau,\boldsymbol{x}))\frac{\partial\boldsymbol{\phi}^{n}_{\rm b}(\tau,\boldsymbol{x})}{\partial\boldsymbol{x}}\dot{\boldsymbol{x}}$ $\geq-\alpha(h(\boldsymbol{\phi}^{n}_{\rm b}(\tau,\boldsymbol{x}))),$ (10a)
$\!\!\nabla h_{\rm b}(\boldsymbol{\phi}^{n}_{\rm b}(T,\boldsymbol{x}))\frac{\partial\boldsymbol{\phi}^{n}_{\rm b}(T,\boldsymbol{x})}{\partial\boldsymbol{x}}\dot{\boldsymbol{x}}$ $\geq-\alpha_{\rm b}(h_{\rm b}(\boldsymbol{\phi}^{n}_{\rm b}(T,\boldsymbol{x}))),$ (10b)
for all ${\tau\in[0,T]}$, where ${\dot{\boldsymbol{x}}=f(\boldsymbol{x})+g(\boldsymbol{x})\boldsymbol{u}}$. Here, ${\boldsymbol{\Phi}_{\rm b}^{n}(\tau,\boldsymbol{x})\triangleq\partial\boldsymbol{\phi}^{n}_{\rm b}(\tau,\boldsymbol{x})/\partial\boldsymbol{x}}$ is the sensitivity matrix, or state-transition matrix (STM), which captures the sensitivity of the flow to perturbations in the initial condition $\boldsymbol{x}$. The sensitivity matrix is the solution to the differential equation
$\begin{gathered}{\dot{\boldsymbol{\Phi}}}_{\rm b}^{n}(\tau,\boldsymbol{x})=F_{\rm cl}(\boldsymbol{\phi}^{n}_{\rm b}(\tau,\boldsymbol{x}))\boldsymbol{\Phi}_{\rm b}^{n}(\tau,\boldsymbol{x}),\\ \mkern 9.0mu\boldsymbol{\Phi}_{\rm b}^{n}(0,\boldsymbol{x})=\boldsymbol{I}_{n\times n},\end{gathered}$ (11)
where $F_{\rm cl}$ is the Jacobian of the closed-loop backup dynamics (4) evaluated at $\boldsymbol{\phi}^{n}_{\rm b}(\tau,\boldsymbol{x})$ and $\boldsymbol{I}_{n\times n}$ is the $n\times n$ identity matrix. This restricts the backup controller to be smooth. While in some special cases a closed-form solution of the STM is known, in general the STM can be numerically integrated along with the backup flow.
The inequality constraints given by (10a) represent an infinite number of constraints. Therefore, to construct a tractable optimization problem they are discretized and enforced at discrete times along the flow. To ensure safety between sample points, (10a) is tightened using a constant $\varepsilon_{\Delta}$:
$\varepsilon_{\Delta}\geq\frac{\Delta}{2}\mathcal{L}_{h}\sup_{\boldsymbol{x}\in\mathcal{C}_{\rm S}}\left\|f_{\rm cl}(\boldsymbol{x})\right\|,$ (12)
where $\Delta\in\mathbb{R}_{>0}$ is a discretization time step satisfying $T/\Delta\in\mathbb{N}$, $\mathcal{L}_{h}\in\mathbb{R}_{>0}$ is the Lipschitz constant of $h$ with respect to the Euclidean norm and $\sup_{\boldsymbol{x}\in\mathcal{C}_{\rm S}}\left\|f_{\rm cl}(\boldsymbol{x})\right\|$ is the maximal velocity of the backup vector field.
As before, the safety of (1) can be enforced for an arbitrary primary controller, $\boldsymbol{u}_{\rm p}\in\mathcal{U}$, by solving the backup CBF optimization problem for the safe control with the newly tightened, implicitly defined constraints:
$\begin{gathered}\boldsymbol{u}_{\rm safe}=\underset{\boldsymbol{u}\in\mathcal{U}}{\text{argmin}}\mkern 9.0mu\frac{1}{2}\left\|\boldsymbol{u}_{\rm p}-\boldsymbol{u}\right\|^{2}\quad\quad\quad(\text{Backup CBF-QP})\\ \text{s.t. }\\ \nabla h(\boldsymbol{\phi}^{n}_{\rm b}(\tau,\boldsymbol{x}))\boldsymbol{\Phi}_{\rm b}^{n}(\tau,\boldsymbol{x})\dot{\boldsymbol{x}}\geq-\alpha(h(\boldsymbol{\phi}^{n}_{\rm b}(\tau,\boldsymbol{x}))-\varepsilon_{\Delta}),\\ \mkern-6.0mu\nabla h_{\rm b}(\boldsymbol{\phi}^{n}_{\rm b}(T,\boldsymbol{x}))\boldsymbol{\Phi}_{\rm b}^{n}(T,\boldsymbol{x})\dot{\boldsymbol{x}}\geq-\alpha_{\rm b}(h_{\rm b}(\boldsymbol{\phi}^{n}_{\rm b}(T,\boldsymbol{x}))),\end{gathered}$
for all $\tau\in\{0,\Delta,\dots,T\}$, where ${\dot{\boldsymbol{x}}=f(\boldsymbol{x})+g(\boldsymbol{x})\boldsymbol{u}}$.
III Disturbance Robustness
While the (Backup CBF-QP) can guarantee safety for a dynamical system in which the dynamics are perfectly known, in practice there will always be unmodeled parameters or external disturbances which will perturb the dynamics. Therefore, it is desirable to leverage the advantages offered by the backup CBF approach, for dynamics with process disturbances. As such, consider the modified dynamical system
$\dot{\boldsymbol{x}}=f(\boldsymbol{x})+g(\boldsymbol{x})\boldsymbol{u}+\boldsymbol{d}_{x},$ (13)
where ${\boldsymbol{d}_{x}\in\mathcal{D}_{x}\subseteq\mathbb{R}^{n}}$ is an unknown additive process disturbance and there exists a constant $\xi\in\mathbb{R}_{>0}$ such that ${\left\|\boldsymbol{d}_{x}\right\|\leq\xi}$. For an initial condition $\boldsymbol{x}(0)=\boldsymbol{x}_{0}\in\mathcal{X}$ and a locally Lipschitz controller $\boldsymbol{u}=k(\boldsymbol{x})$, if $\boldsymbol{d}_{x}$ is piecewise continuous in time, the closed-loop system (13) has a unique solution $\boldsymbol{\phi}^{d}(t,\boldsymbol{x}_{0})$ over an interval of existence.
We assume that a backup set $\mathcal{C}_{\rm B}$ and corresponding control law $\boldsymbol{u}_{\rm b}$ can be obtained which render a small subset of $\mathcal{C}_{\rm S}$ robustly forward invariant. This is made more precise below.
Assumption 1.
The backup controller $\boldsymbol{u}_{\rm b}$ renders the backup set $\mathcal{C}_{\rm B}$ forward invariant along (13) for any disturbance $\boldsymbol{d}_{x}$ which satisfies $\left\|\boldsymbol{d}_{x}\right\|\leq\xi$.
Next we define two separate flows: the nominal and the disturbed backup flow. The nominal backup flow $\boldsymbol{\phi}^{n}_{\rm b}(\tau,\boldsymbol{x})$ satisfies (5) under the robust control law $\boldsymbol{u}_{\rm b}$, while the disturbed backup flow, denoted $\boldsymbol{\phi}^{d}_{\rm b}(\tau,\boldsymbol{x})$, is the solution to
${\dot{\boldsymbol{\phi}^{d}_{\rm b}\mkern-1.0mu}\mkern 1.0mu}{}(\tau,\boldsymbol{x})=f_{\rm cl}(\boldsymbol{\phi}^{d}_{\rm b}(\tau,\boldsymbol{x}))+\boldsymbol{d}_{x},\mkern 9.0mu\boldsymbol{\phi}^{d}_{\rm b}(0,\boldsymbol{x})=\boldsymbol{x}.$ (14)
Again, it is assumed that for any $\boldsymbol{x}\in\mathcal{X}$ there exists a unique solution $\boldsymbol{\phi}^{d}_{\rm b}:[0,T]\times\mathcal{X}\rightarrow\mathcal{X}$ to (14). Consider $\mathcal{C}_{\rm D}\subseteq\mathcal{C}_{\rm S}$
$\mathcal{C}_{\rm D}\triangleq\left\{\boldsymbol{x}\in\mathcal{X}\,\middle|\,\begin{array}[]{c}h(\boldsymbol{\phi}^{d}_{\rm b}(\tau,\boldsymbol{x}))\geq 0,\forall\mkern 1.0mu\tau\in[0,T],\\ h_{\rm b}(\boldsymbol{\phi}^{d}_{\rm b}(T,\boldsymbol{x}))\geq 0\\ \end{array}\right\}.$ (17)
We are interested in forward invariance conditions for $\mathcal{C}_{\rm D}$, but since the disturbance is unknown, we will instead derive forward invariance conditions for a set which over-approximates the disturbed flow. Consider a new set, $\mathcal{C}_{\rm I}$, defined by
$\mathcal{C}_{\rm I}\triangleq\left\{\boldsymbol{x}\in\mathcal{X}\,\middle|\,\begin{array}[]{c}h(\boldsymbol{\phi}^{n}_{\rm b}(\tau,\boldsymbol{x}))\geq\epsilon_{\tau},\forall\mkern 1.0mu\tau\in[0,T],\\ h_{\rm b}(\boldsymbol{\phi}^{n}_{\rm b}(T,\boldsymbol{x}))\geq\epsilon_{\rm b}\\ \end{array}\right\}.$ (20)
The set is entirely governed by the nominal trajectory and additional tightening terms $\epsilon_{\tau}$ and $\epsilon_{\rm b}$. For judiciously chosen values of $\epsilon_{\tau}$ and $\epsilon_{\rm b}$, we show that $\mathcal{C}_{\rm I}$ is a subset of $\mathcal{C}_{\rm D}$.
Lemma 1.
Let $\mathcal{L}_{h}$ and $\mathcal{L}_{h_{\rm b}}$ be the Lipschitz constants of $h$ and $h_{\rm b}$ respectively, and let $\delta_{{\rm max}}(\tau)$ be a norm bound on the deviation between $\boldsymbol{\phi}^{n}_{\rm b}(\tau,\boldsymbol{x})$ and $\boldsymbol{\phi}^{d}_{\rm b}(\tau,\boldsymbol{x})$ at time $\tau\in[0,T]$:
$\|\boldsymbol{\phi}^{n}_{\rm b}(\tau,\boldsymbol{x})-\boldsymbol{\phi}^{d}_{\rm b}(\tau,\boldsymbol{x})\|\leq\delta_{{\rm max}}(\tau),$ (21)
for all $\boldsymbol{x}\in\mathcal{C}_{\rm S}$. If $\epsilon_{\tau}\geq\mathcal{L}_{h}\delta_{{\rm max}}(\tau)$ holds for all $\tau\in[0,T]$ and $\epsilon_{\rm b}\geq\mathcal{L}_{h_{\rm b}}\delta_{{\rm max}}(T)$ also holds, then $\mathcal{C}_{\rm I}\subseteq\mathcal{C}_{\rm D}$.
Proof.
Consider any state $\boldsymbol{x}\in\mathcal{C}_{\rm I}$. Membership to $\mathcal{C}_{\rm I}$ implies that $h(\boldsymbol{\phi}^{n}_{\rm b}(\tau,\boldsymbol{x}))\geq\epsilon_{\tau}\geq\mathcal{L}_{h}\delta_{{\rm max}}(\tau)$. Hence it follows that
$h(\boldsymbol{\phi}^{d}_{\rm b}(\tau,\boldsymbol{x}))$ $=h(\boldsymbol{\phi}^{n}_{\rm b}(\tau,\boldsymbol{x}))-\big{(}h(\boldsymbol{\phi}^{n}_{\rm b}(\tau,\boldsymbol{x}))-h(\boldsymbol{\phi}^{d}_{\rm b}(\tau,\boldsymbol{x}))\big{)}$
$\geq\mathcal{L}_{h}\delta_{{\rm max}}(\tau)-\big{|}h(\boldsymbol{\phi}^{n}_{\rm b}(\tau,\boldsymbol{x}))-h(\boldsymbol{\phi}^{d}_{\rm b}(\tau,\boldsymbol{x}))\big{|}.$
By Lipschitz continuity of the constraint function $h$
$|h(\boldsymbol{\phi}^{n}_{\rm b}(\tau,\boldsymbol{x}))-h(\boldsymbol{\phi}^{d}_{\rm b}(\tau,\boldsymbol{x}))|$
$\quad\leq\mathcal{L}_{h}\left\|\boldsymbol{\phi}^{n}_{\rm b}(\tau,\boldsymbol{x})-\boldsymbol{\phi}^{d}_{\rm b}(\tau,\boldsymbol{x})\right\|\leq\mathcal{L}_{h}\delta_{{\rm max}}(\tau),$
we obtain ${h(\boldsymbol{\phi}^{d}_{\rm b}(\tau,\boldsymbol{x}))\geq 0}$ for any $\boldsymbol{x}\in\mathcal{C}_{\rm I}$. Similar logic can be applied to the constraint on the reachability of the backup set. For any $\boldsymbol{x}\in\mathcal{C}_{\rm I}$, the nominal backup trajectory from $\boldsymbol{x}$ satisfies $h_{\rm b}(\boldsymbol{\phi}^{n}_{\rm b}(T,\boldsymbol{x}))\geq\epsilon_{\rm b}\geq\mathcal{L}_{h_{\rm b}}\delta_{{\rm max}}(T)$, and we have
$|h_{\rm b}(\boldsymbol{\phi}^{n}_{\rm b}(T,\boldsymbol{x}))-h_{\rm b}(\boldsymbol{\phi}^{d}_{\rm b}(T,\boldsymbol{x}))|\leq\mathcal{L}_{h_{\rm b}}\delta_{{\rm max}}(T).$
These guarantee that $h_{\rm b}(\boldsymbol{\phi}^{d}_{\rm b}(T,\boldsymbol{x}))\geq 0$. Thus, all the functions which define $\mathcal{C}_{\rm D}$ are positive, meaning that $\boldsymbol{x}\in\mathcal{C}_{\rm D}$ for all $\boldsymbol{x}\in\mathcal{C}_{\rm I}$, and so $\mathcal{C}_{\rm I}\subseteq\mathcal{C}_{\rm D}$. ∎
Lemma 1 assumes that a time-varying bound on the deviation between the nominal and disturbed backup flow, $\delta_{\rm max}(\tau)$, can be found. While problem-specific bounds can be obtained, we utilize a generalization of the Gronwall-Bellman inequality to obtain such a bound for a wide class of nonlinear systems.
Lemma 2 (Theorem 2.5 in [12]).
For systems (5) and (14), let $f_{\rm cl}$ be locally Lipschitz on $\mathcal{X}$ with Lipschitz constant $\mathcal{L}_{\rm cl}$ and $\boldsymbol{d}_{x}$ be piecewise continuous in $\tau$ on $[0,T]$. If $\left\|\boldsymbol{d}_{x}\right\|\leq\xi$ for all $\boldsymbol{d}_{x}$ and some $\xi>0$ then
$\left\|\boldsymbol{\phi}^{n}_{\rm b}(\tau,\boldsymbol{x})-\boldsymbol{\phi}^{d}_{\rm b}(\tau,\boldsymbol{x})\right\|\leq\frac{\xi}{\mathcal{L}_{\rm cl}}(e^{\mathcal{L}_{\rm cl}\tau}-1)\triangleq\delta_{{\rm max}}(\tau),$
for all $\tau\in[0,T]$.
Remark 1.
In the case where contraction bounds on the deviation between the disturbed and nominal backup flow can be obtained, $\delta_{{\rm max}}(\tau)$ can be made much less conservative. Details on such bounds can be found in [13, Corollary 3.17]. A contraction bound is used effectively in the spacecraft rotation example in Section IV-B. For linear systems, flow deviation bounds can be made even tighter.
Using the definition of $\mathcal{C}_{\rm D}$ and the corresponding robust backup controller, we now examine the properties of $\mathcal{C}_{\rm D}$.
Lemma 3.
The set $\mathcal{C}_{\rm D}$ is controlled invariant, and the robust backup controller $\boldsymbol{u}_{\rm b}$ renders $\mathcal{C}_{\rm D}$ forward invariant along (13), such that
$\boldsymbol{x}\in\mathcal{C}_{\rm D}\implies\boldsymbol{\phi}_{\rm b}^{d}(\vartheta,\boldsymbol{x})\in\mathcal{C}_{\rm D},\forall\mkern 1.0mu\vartheta\geq 0.$ (22)
Proof.
From the definition of $\mathcal{C}_{\rm D}$ and with 1
$\boldsymbol{x}\in\mathcal{C}_{\rm D}\implies\boldsymbol{\phi}_{\rm b}^{d}(\tau,\boldsymbol{x})\in\mathcal{C}_{\rm B}\subseteq\mathcal{C}_{\rm S},\forall\mkern 1.0mu\tau\geq T.$ (23)
By definition, the flow is recursive in nature and thus for any $\boldsymbol{x}\in\mathbb{R}^{n}$ and $\tau,\vartheta\geq 0$, $\boldsymbol{\phi}_{\rm b}^{d}(\tau+\vartheta,\boldsymbol{x})=\boldsymbol{\phi}_{\rm b}^{d}(\tau,\boldsymbol{\phi}_{\rm b}^{d}(\vartheta,\boldsymbol{x}))$. Using (23) and the recursive property of the flow
$\boldsymbol{x}\in\mathcal{C}_{\rm D}\implies\boldsymbol{\phi}_{\rm b}^{d}(T,\boldsymbol{\phi}_{\rm b}^{d}(\vartheta,\boldsymbol{x}))\in\mathcal{C}_{\rm B},\forall\mkern 1.0mu\vartheta\geq 0.$ (24)
From (23) and by definition (17) $\boldsymbol{x}\in\mathcal{C}_{\rm D}\implies\boldsymbol{\phi}_{\rm b}^{d}(\tau,\boldsymbol{x})\in\mathcal{C}_{\rm S},\forall\mkern 1.0mu\tau\geq 0$. Using the recursive property once more
$\mkern-3.0mu\boldsymbol{x}\in\mathcal{C}_{\rm D}\mkern-6.0mu\implies\mkern-6.0mu\boldsymbol{\phi}_{\rm b}^{d}(\tau,\boldsymbol{\phi}_{\rm b}^{d}(\vartheta,\boldsymbol{x}))\mkern-2.0mu\in\mkern-2.0mu\mathcal{C}_{\rm S},\forall\mkern 1.0mu\tau\mkern-1.0mu\in\mkern-1.0mu[0,T],\forall\vartheta\mkern-2.0mu\geq\mkern-2.0mu0.$ (25)
Therefore applying definition (17) with (24) and (25) completes the proof. ∎
While the controlled invariance of $\mathcal{C}_{\rm D}$ has been established, the conditions on $\boldsymbol{u}$ for forward invariance cannot yet be obtained as $\mathcal{C}_{\rm D}$ itself is unknown. This motivates the following theorems.
Theorem 2.
For any $\boldsymbol{x}\in\mathcal{C}_{\rm I}$, there exists a controller $\boldsymbol{u}$ such that $\boldsymbol{\phi}^{d}(\vartheta,\boldsymbol{x})\in\mathcal{C}_{\rm D}\subseteq\mathcal{C}_{\rm S},\forall\mkern 2.0mu\vartheta\geq 0$.
Proof.
By Lemma 1, any $\boldsymbol{x}\mkern-3.0mu\in\mkern-2.0mu\mathcal{C}_{\rm I}\mkern-6.0mu\implies\mkern-6.0mu\boldsymbol{x}\mkern-3.0mu\in\mkern-3.0mu\mathcal{C}_{\rm D}$ and by Lemma 3, $\boldsymbol{u}_{\rm b}$ is one such controller which ensures that $\boldsymbol{\phi}^{d}(\vartheta,\boldsymbol{x})\in\mathcal{C}_{\rm D}\subseteq\mathcal{C}_{\rm S},\forall\mkern 2.0mu\vartheta\geq 0$. ∎
We are now ready to apply a safety condition to obtain conditions for a controller which ensures safety of (13). From the definition of $\mathcal{C}_{\rm I}$ we have
$\begin{gathered}\dot{h}(\boldsymbol{\phi}^{n}_{\rm b}(\tau,\boldsymbol{x}),\boldsymbol{u})\geq-\alpha(h(\boldsymbol{\phi}^{n}_{\rm b}(\tau,\boldsymbol{x}))-\epsilon_{\tau}),\\ \dot{h}_{\rm b}(\boldsymbol{\phi}^{n}_{\rm b}(T,\boldsymbol{x}),\boldsymbol{u})\geq-\alpha_{\rm b}(h_{\rm b}(\boldsymbol{\phi}^{n}_{\rm b}(T,\boldsymbol{x}))-\epsilon_{\rm b}),\end{gathered}$ (28)
where by expanding the total derivatives for system (13) this becomes, $\forall\tau\in[0,T],$
$\begin{gathered}\mkern-4.0mu\nabla h(\boldsymbol{\phi}^{n}_{\rm b}(\tau,\boldsymbol{x}))\boldsymbol{\Phi}_{\rm b}^{n}(\tau,\boldsymbol{x})\dot{\boldsymbol{x}}^{d}\geq-\alpha(h(\boldsymbol{\phi}^{n}_{\rm b}(\tau,\boldsymbol{x}))-\epsilon_{\tau}),\\ \nabla h_{\rm b}(\boldsymbol{\phi}^{n}_{\rm b}(T,\boldsymbol{x}))\boldsymbol{\Phi}_{\rm b}^{n}(T,\boldsymbol{x})\dot{\boldsymbol{x}}^{d}\geq-\alpha_{\rm b}(h_{\rm b}(\boldsymbol{\phi}^{n}_{\rm b}(T,\boldsymbol{x}))-\epsilon_{\rm b}).\end{gathered}$ (31)
Here, $\dot{\boldsymbol{x}}^{d}\triangleq f(\boldsymbol{x})+g(\boldsymbol{x})\boldsymbol{u}+\boldsymbol{d}_{x}$. Using this expansion, we can show that a control solution which realizes forward invariance of $\mathcal{C}_{\rm I}$ keeps the disturbed system safe, and that conditions for such a controller can be directly computed, despite the unknown disturbance.
Theorem 3.
If any controller $\boldsymbol{u}$ satisfies
$\begin{gathered}\nabla h(\boldsymbol{\phi}^{n}_{\rm b}(\tau,\boldsymbol{x}))\boldsymbol{\Phi}_{\rm b}^{n}(\tau,\boldsymbol{x})\big{(}f(\boldsymbol{x})+g(\boldsymbol{x})\boldsymbol{u}\big{)}-\eta\geq\\ \hskip 56.9055pt-\alpha(h(\boldsymbol{\phi}^{n}_{\rm b}(\tau,\boldsymbol{x}))-\epsilon_{\tau}),\\ \nabla h_{\rm b}(\boldsymbol{\phi}^{n}_{\rm b}(T,\boldsymbol{x}))\boldsymbol{\Phi}_{\rm b}^{n}(T,\boldsymbol{x})\big{(}f(\boldsymbol{x})+g(\boldsymbol{x})\boldsymbol{u}\big{)}-\eta_{\rm b}\geq\\ \hskip 68.28644pt-\alpha_{\rm b}(h_{\rm b}(\boldsymbol{\phi}^{n}_{\rm b}(T,\boldsymbol{x}))-\epsilon_{\rm b}),\end{gathered}$ (36)
with robustness terms defined by
$\begin{gathered}\eta\triangleq\xi\left\|\nabla h(\boldsymbol{\phi}^{n}_{\rm b}(\tau,\boldsymbol{x}))\boldsymbol{\Phi}_{\rm b}^{n}(\tau,\boldsymbol{x})\right\|,\\ \eta_{\rm b}\triangleq\xi\left\|\nabla h_{\rm b}(\boldsymbol{\phi}^{n}_{\rm b}(T,\boldsymbol{x}))\boldsymbol{\Phi}_{\rm b}^{n}(T,\boldsymbol{x})\right\|,\end{gathered}$
then $\boldsymbol{x}_{0}\in\mathcal{C}_{\rm I}\implies\boldsymbol{\phi}^{d}(t,\boldsymbol{x}_{0})\in\mathcal{C}_{\rm I}\subseteq\mathcal{C}_{\rm D}\subseteq\mathcal{C}_{\rm S},$ for all $\mkern 1.0mut>0$.
Proof.
From a direct application of Theorem 1 for dynamical system (13), we obtain (31). As done in [5], the robustness terms $\eta$ and $\eta_{\rm b}$ upperbound the unknown $\boldsymbol{d}_{x}$ term. The condition (36) thus ensures that for any $\boldsymbol{x}_{0}\in\mathcal{C}_{\rm I}$, $\boldsymbol{\phi}^{d}(t,\boldsymbol{x}_{0})\in\mathcal{C}_{\rm I},\forall\mkern 1.0mut>0,$ and from Lemma 1, $\mathcal{C}_{\rm I}\subseteq\mathcal{C}_{\rm D}$. ∎
Naturally, the original backup CBF constraints (10) are recovered in the absence of disturbances (i.e., $\xi=0$). As the constraints in (36) are continuous in $\tau$, the trajectory is again discretized and appropriately tightened via a constant term $\varepsilon_{\Delta}$ where
$\varepsilon_{\Delta}\geq\frac{\Delta}{2}\mathcal{L}_{h}(\sup_{\boldsymbol{x}\in\mathcal{C}_{\rm S}}\left\|f_{\rm cl}(\boldsymbol{x})\right\|+\xi).$ (37)
The result of Theorem 3 is now ready to be directly utilized in a new point-wise optimal controller accounting for disturbances. The Disturbance-Robust Backup CBF (DR-BCBF) optimization problem is written as:
$\begin{gathered}\boldsymbol{u}_{\rm safe}=\underset{\boldsymbol{u}\in\mathcal{U}}{\text{argmin}}\mkern 9.0mu\frac{1}{2}\left\|\boldsymbol{u}_{\rm p}-\boldsymbol{u}\right\|^{2}\quad\quad\quad\quad(\text{DR-BCBF-QP})\\ \text{s.t. }\\ \nabla h(\boldsymbol{\phi}^{n}_{\rm b}(\tau,\boldsymbol{x}))\boldsymbol{\Phi}_{\rm b}^{n}(\tau,\boldsymbol{x})\big{(}f(\boldsymbol{x})+g(\boldsymbol{x})\boldsymbol{u}\big{)}-\eta\geq\\ \hskip 80.00012pt\ignorespaces-\alpha(h(\boldsymbol{\phi}^{n}_{\rm b}(\tau,\boldsymbol{x}))-\epsilon_{\tau}-\varepsilon_{\Delta}),\\ \nabla h_{\rm b}(\boldsymbol{\phi}^{n}_{\rm b}(T,\boldsymbol{x}))\boldsymbol{\Phi}_{\rm b}^{n}(T,\boldsymbol{x})\big{(}f(\boldsymbol{x})+g(\boldsymbol{x})\boldsymbol{u}\big{)}-\eta_{\rm b}\geq\\ \hskip 90.00014pt\ignorespaces-\alpha_{\rm b}(h_{\rm b}(\boldsymbol{\phi}^{n}_{\rm b}(T,\boldsymbol{x}))-\epsilon_{\rm b}),\end{gathered}$
for all $\tau\in\{0,\Delta,\dots,T\}$.
Remark 2.
From Theorem 2, $\mathcal{C}_{\rm D}$ is controlled invariant, however the controlled invariance of $\mathcal{C}_{\rm I}$ itself cannot be proven without additional assumptions on $\boldsymbol{u}_{\rm b}$ and $\mathcal{C}_{\rm B}$. Therefore, feasibility of the optimization problem (DR-BCBF-QP) is not guaranteed. However, in the case that the optimization problem becomes infeasible, the robust backup control law $\boldsymbol{u}_{\rm b}$ can be used to stay in $\mathcal{C}_{\rm D}$ until the optimization problem becomes feasible again.
IV Numerical Results
To demonstrate the effectiveness of the proposed method at assuring safety under bounded disturbances, the results from two numerical examples are shown: a double integrator and a rigid body spacecraft with angular rate constraints.
IV-A Double Integrator
Consider a simple example of a double integrator given by
$\dot{\boldsymbol{x}}=\begin{bmatrix}x_{2}\\ u\end{bmatrix}+\boldsymbol{d}_{x},$ (38)
with a state vector $\boldsymbol{x}=[x_{1},x_{2}]^{T}\in\mathbb{R}^{2}$ where $x_{1}$ is the position and $x_{2}$ is the velocity, and an acceleration control variable $u\in\mathcal{U}=[-1,1]$. The safe set is defined as ${\mathcal{C}_{\rm S}\triangleq\{\boldsymbol{x}\in\mathbb{R}^{2}:-x_{1}\geq 0\}}$. $\boldsymbol{d}_{x}\in\mathcal{D}_{x}\subseteq\mathbb{R}^{n}$ is an unknown additive process disturbance and there exists a constant $\xi\in\mathbb{R}_{>0}$ such that $\left\|\boldsymbol{d}_{x}\right\|\leq\xi$. The backup control law $\boldsymbol{u}_{\rm b}(\boldsymbol{x})=-1$ will bring the system to the backup set ${\mathcal{C}_{\rm B}\triangleq\{\boldsymbol{x}\in\mathbb{R}^{2}:-x_{1}\geq 0,-x_{2}\geq 0\}}$ as long as $\xi<1$. In this particular example, the unknown disturbance is constant, with ${\boldsymbol{d}_{x}=\xi\frac{\boldsymbol{v}}{\left\|\boldsymbol{v}\right\|}}$ where ${\boldsymbol{v}=[1,1]^{T}}$, and ${\xi=0.08}$. The discretization time step used for computing the nominal backup flow is ${\Delta=0.02\mkern 4.0mu{\rm s}}$. The primary control law is ${\boldsymbol{u}_{\rm p}=1}$ which drives the system to the right half-plane (unsafe region).
Using the proposed disturbance-robust backup CBF approach, the forward invariant set $\mathcal{C}_{\rm I}$ is computed for various values of $T$ between $0.5\mkern 4.0mu{\rm s}$ and $1.25\mkern 4.0mu{\rm s}$. These sets are plotted on Figure 2 in pink. The backup set, $\mathcal{C}_{\rm B}$, is plotted in blue and a robust controlled invariant set $\mathcal{C}_{\rm R}$ is plotted in green. For this simple linear system such a set can be computed by analytically solving for the flow and accounting for worst-case disturbances. The black dotted line represents the trajectory of the disturbed system using the proposed controller, with an integration horizon $T=1.25\mkern 4.0mu{\rm s}$. Notably, safety is maintained along this trajectory, as stated by Theorem 3. Nominal backup trajectories in gray emanate from the disturbed trajectory and the gray circles centered around them are the Gronwall norm balls from Lemma 2. The final Gronwall norm ball, colored in red, is always contained in $\mathcal{C}_{\rm B}$, as required by (20).
The plots show that as the backup horizon $T$ is increased, the forward invariant set $\mathcal{C}_{\rm I}$ increases in size, up to a certain point. Since the Gronwall bound grows exponentially with time, the longer the backup horizon, the larger the final bound, $\delta_{\rm max}(T)$, will be. If $T$ is too large, the constraint on the terminal point of the nominal backup trajectory will dominate, and the set $\mathcal{C}_{\rm I}$ will begin to shrink. This therefore introduces a trade-off as $T$ cannot be made arbitrarily large when using Lemma 2. Naturally, as the disturbance bound $\xi$ increases, the size of $\mathcal{C}_{\rm I}$ will shrink since the bound on the backup flows is proportional to $\xi$.
Figure 2: Phase space visualization of safety-critical control of double integrator system under bounded disturbances, using the proposed disturbance-robust backup control barrier function approach.
IV-B Rigid Body Spacecraft Rotation
Figure 3: Safety-critical control of rigid body spacecraft using the proposed disturbance-robust backup control barrier function approach. (Left) State-space visualization of angular velocity components showing the trajectory of the angular velocity vector over time. (Right) Angular velocity norm over time showing adherence to safety constraint (top), and commanded primary control and actual (safe) control signal over time (bottom).
Consider next an example of a rigid body spacecraft with a known inertia tensor in the body frame given by $\boldsymbol{J}$ where the dynamics of the angular velocities can be described by Euler’s rotational equations of motion
$\boldsymbol{\dot{\omega}}=\boldsymbol{J}^{-1}(-\boldsymbol{\omega}\times\boldsymbol{J}\boldsymbol{\omega}+\boldsymbol{u})+\boldsymbol{d}_{x}.$ (39)
Here, $\boldsymbol{u}$ is the control torque vector that can be applied by the spacecraft to control the angular velocity and $\boldsymbol{d}_{x}$ is an unknown but bounded additive disturbance vector, such that $\left\|\boldsymbol{d}_{x}\right\|\leq\xi\in\mathbb{R}_{>0}$. The states of interest are the angular velocity vector elements in the body frame, $\boldsymbol{\omega}=[\omega_{1},\omega_{2},\omega_{3}]^{T}$.
The safety objective is to ensure that the norm of the angular velocity vector of the spacecraft does not exceed a maximum value. This constraint is designed to prevent damage to sensors that may be onboard the rigid body spacecraft. The safe set is therefore $\mathcal{C}_{\rm S}=\{\boldsymbol{\omega}\in\mathbb{R}^{3}:h(\boldsymbol{\omega})\geq 0\}$ where $h(\boldsymbol{\omega})=\omega_{\rm max}^{2}-\left\|\boldsymbol{\omega}\right\|^{2}\geq 0$ and $\omega_{\rm max}\in\mathbb{R}_{>0}$ represents the maximum allowable angular velocity. The primary controller is given by ${\boldsymbol{u}_{\rm p}(t)=\text{sin}([\frac{t}{2},\frac{t}{2}-\frac{\pi}{4},\frac{t}{4}+\frac{\pi}{4}]^{T})}$. The robust backup control law $\boldsymbol{u}_{\rm b}(\boldsymbol{\omega})=-k_{b}\boldsymbol{J}\boldsymbol{\omega}+\boldsymbol{\omega}\times\boldsymbol{J}\boldsymbol{\omega}$ renders the backup set $\mathcal{C}_{\rm B}\triangleq\{\boldsymbol{\omega}\in\mathbb{R}^{3}:h_{\rm b}(\boldsymbol{\omega})=\gamma-\frac{1}{2}\boldsymbol{\omega}^{T}\boldsymbol{J}\boldsymbol{\omega}\geq 0\}$ robustly forward invariant for sufficiently large gain $k_{b}$. $\mathcal{C}_{\rm B}$ is a level set of the spacecraft’s rotational energy, defined by the scalar $\gamma$. It can be shown that $k_{b}>(\lambda_{\rm max}\xi)/(\sqrt{2\gamma\lambda_{\rm min}})$ ensures $\dot{h}_{\rm b}(\boldsymbol{\omega},\boldsymbol{u})\geq 0$ at the boundary of the level set for (39), where $\lambda_{\rm max}$ and $\lambda_{\rm min}$ are the maximum and minimum eigenvalues of $\boldsymbol{J}$, respectively. The proof is omitted for brevity.
It is straightforward to verify that the closed-loop nominal backup dynamics are strongly contracting with a rate of $k_{b}$ since $f_{\rm cl}(\boldsymbol{\omega})=-k_{b}\boldsymbol{\omega}$. Because the log norm of the closed-loop Jacobian ($\partial f_{\rm cl}/\partial\boldsymbol{\omega}$) is upper bounded by $-k_{b}$, the disturbed and nominal backup flows can be bounded as
$\left\|\boldsymbol{\phi}^{n}_{\rm b}(\tau,\boldsymbol{\omega})-\boldsymbol{\phi}^{d}_{\rm b}(\tau,\boldsymbol{\omega})\right\|\leq\frac{\xi}{k_{b}}(1-e^{-k_{b}\tau}),$ (40)
by [13, Corollary 3.17]. This yields much tighter $\epsilon_{\tau}$ and $\epsilon_{\rm b}$ terms than the general Gronwall bound.
For the simulations, $\omega_{\rm max}=1\mkern 4.0mu\rm rad/s$, $\gamma=2\rm\mkern 4.0muJ$, $\xi=0.1\mkern 4.0mu{\rm rad/s^{2}}$ and $\boldsymbol{J}=\texttt{diag}\{[12,12,5]\}\mkern 4.0mu{\rm kg\mkern 1.0mum^{2}}$. The $\texttt{diag}\{\cdot\}$ operator returns a diagonal matrix with the inputted vector on the diagonal. The discretization time step used for computing the nominal backup flow was $\Delta=0.05\mkern 4.0mu{\rm s}$ and the integration horizon was $T=1.75\mkern 4.0mu{\rm s}$. The disturbance vector was time-varying, given by $\boldsymbol{d}_{x}(t)=\xi\frac{\boldsymbol{v}(t)}{\|\boldsymbol{v}(t)\|}$ where $\boldsymbol{v}(t)=\text{sin}([\frac{t}{2}+\frac{\pi}{2},\frac{t}{2},\frac{t}{2}-\frac{\pi}{2}]^{T})$.
Figure 3 shows the results of using the disturbance-robust backup CBF method for this application with the contraction bounds in (40). On the left of Figure 3 the trajectory of the disturbed system over time is plotted in blue, and the goal is to keep this trajectory within the red sphere. The magenta sections of the trajectory indicate that the primary control signal, $\boldsymbol{u}_{\rm p}$, has been modified to assure safety. The wire-frame spheres represent the contraction norm balls along the nominal backup flow in cyan. Note that as required by (20), the terminal spheres are entirely contained within the $\mathcal{C}_{\rm B}$ ellipse. The right plots of Figure 3 show that the norm constraint on the angular velocity is adhered to in the presence of time-varying disturbances and that the primary control signal is modified only when $\left\|\boldsymbol{\omega}\right\|$ is near $\omega_{\rm max}$.
V Conclusions
This paper presents a novel safety-critical control framework to handle unknown bounded disturbances for a broad class of nonlinear systems. We extend the method of backup CBFs to handle such disturbances by providing forward invariance conditions for a subset of a controlled invariant set governed by the disturbed system. We prove that enforcing these conditions guarantees safety of the disturbed system and demonstrate the effectiveness of the approach with two numerical simulation examples.